# Inferring Capabilities from Task Performance with Bayesian Triangulation

### v1

@@ Appendix E Prediction Results
Table 8: Brier Scores (BS) and the breakdown into calibration (Cal.) and refinement (Ref.) for the Measurement Layouts (ML) and aggregate predictions (Agg.)
Appendix F Methodology for Designing Measurement Layouts
For complex, multifaceted tasks measurement layouts have the potential to grow and become large and complex themselves. It’s therefore important that to apply them to other domains (and reap the benefits of the additional evaluation information they provide) to have a general methodology for constructing the appropriate measurement layout structure. To this end, we developed a methodology for constructing measurement layouts that we wish to explicate here more clearly.
As stated in the main paper, the core of our approach is to identify the demands of a task, and to compare these against the capabilities of the system being evaluated. Tasks are composed of multiple instances that vary in the degree of demands. In our methodology we tie demands with meta-features in order to give the demands values that are tied to objective properties of the task. For each demand in a task, a system will have varying levels of performance at dealing with the demand based on its capability. we model this as a probability of the system successfully handling the aspect of the task with its meta-feature value. This is where the linking function comes into play. We often leveraged the sigmoid function $\sigma(capabiltiy-demand)$. This approach is beneficial because it ties capabilities to probabilistic performance based on the demand. This also causes the capabilities to have the same unit as the demand, and allows us to interpret a capability with value $x$ as consistently succeeding on the demand with meta-feature $x$ in $50\%$ of instances.
Certain demands also require that other demands have been met. This requires a dependency understanding of the demands of the task. Performances for dependent task demands can be incorporated into the sigmoid margin calculation or can have an effect on the resulting performance value. The exact approach for combining demand performances will depend on the task itself, and whether the demands are compensatory—whether performance can be achieved with one capability compensating for lacking another. Our environments were non-compensatory, so we often opted for taking the product of probabilities. However, more complex relations such as taking the minimum or maximum of a set of probabilities is available by utilising a generalised mean. The other major aspect that we need to identify to construct the measurement layout is the sources of possible bias. In principle, this is similar to creating the structure for capabilities: We still identify meta-features that may affect performance, however, now we are looking for those that are not necessary for the task. We still include a bias ”performance” for each instance (how present was the bias in this instance, and in which direction was it directed?) , but this is now included in the sigmoid function for the relevant capability it may affect: $\sigma(capability-demand+bias)$. Once all of the demands and biases are accounted for, a final combination can give an output for overall task performance.
Breaking a complex task down into demands and possible biases is not always straightforward. Care, and domain knowledge of what constitutes success in the task, needs to be utilised to identify the core broad capabilities required, and how these break down to identifiable meta-features. We recommend an iterative approach: If a particular measurement layout cannot explain system behaviour reliably (with a low standard deviation on the resulting posterior distributions), then there is either too few evaluation instances, or the measurement layout is missing an important capability or bias.
When evaluating Artificial Intelligence systems, and for that matter non-human Biological Intelligences, we need to be careful to avoid anthropomorphising the system. We do not want to mistakenly ascribe capabilities or explain behaviours based on a crude facsimile of human behaviour. AI systems, in particular, have the potential to solve tasks using incredibly different internal methods and skills. We have been careful to avoid this mistake in our work by focusing on the required demands for a task rather than any model for how the system internally processes meta-features.
Appendix G Computation Code, and Datasets
@@
Our experiments did not require a large amount of compute, We ran all of our experiments on Google Colaboratory (Google 2017) using the free tier of service without utilising the GPU. All corresponding code, datasets and raw results will be released on publication to ensure ease of replication.

### v2

@@ E.3 Object Permanence Predictive Results
Table 9: Brier Scores (BS) and the breakdown into calibration (Cal.) and refinement (Ref.) for the Measurement Layouts (ML) and aggregate predictions (Agg.). Bold indicates whether the Measurement Layout or Aggregate Brier score was lower (and therefore better).
Appendix F Supplementary Experiment: Animal AI Olympics
This experiment exemplifies that evaluation suites that weren’t designed for this type of nuanced analysis can still provide considerable insights.
The Animal AI Olympics (AAIO) was an open competition evaluating AI agents in the AAI 3D environment, where agents had to get a reward by locating it and navigating in an arena. In the Animal-AI Olympics (Crosby et al., 2020), an example arena containing the possible objects was provided to contestants. They were instructed that their agents would be evaluated on a variety of problems constructed from the objects in the example arena. The contestants then needed to build an agent—using any method of their choosing—that they believed would succeed at a wide variety of tasks. The tasks that ultimately made up the evaluation suite were based on adaptations of tasks from comparative and developmental literature. More information can be found at http://animalai.org/animal-ai-olympics. For our work with the AAIO dataset we used a subset of 69 tasks instances that were solvable with simple goal-directed behaviour. The exact instances can be found in Appendix F.1. We include 68 participant agents for which we know nothing about their design.
F.1 Animal AI Olympics Instances
For our work with the AAIO dataset we used a subset of tasks instances that were solvable with simple goal-directed behaviour. The specific task instances we used were:
The types of arenas that these instances correspond to can be found at (http://animalai.org/animal-ai-olympics).
Dimension Description Range/Values
Reward Size The size of the reward (the sign is modified so lower is larger) $[0,1.9]$
Reward Distance Euclidean distance from the agent’s start position to the reward $[0,5.3]$
Reward Behind Whether the reward originates behind (1), left or right (0.5) or in front (0) of the agent $\{0,0.5,1\}$
XPos Whether the reward is to the left (1-1), the right (1) or is centred (0) relative to the agent’s start position $\{-1,0,1\}$
Note that these ranges determine the ranges of the capabilities, so that we have units. For instance, if distance to the reward were measured in metres then the navigation capability would have metres as a unit. Let the navigation capability of an agent be, say, 20 metres. Since it is defined with a logistic function of distance, that means that the agent can achieve excellent performance when well below 20, performing at chance around the 20 metre mark, and failing every instance when the reward is well beyond 20 metres away.
Figure 23: Inferred cognitive profile elements, along with the success rate, for the 68 agents in the AAIO. The gradient colour red-yellow-green denotes how high the means are (green is best) while the gradient circles white-grey-black denote how high the standard deviations are (black is highest).
Fig. 23 shows the complete inference results for the 68 agents in the AAIO. The agents are ordered by the aggregate success rate at the collection of task instances. We see that, for the agents in the middle of the table in particular, our approach is able to decompose agent capabilities at this task into the constituent navigation and visual abilities. Bias has a more diverse behaviour throughout the table, with strong biases for extremely low scoring agents too, such as with agents ‘ahorizon’ and ‘Octopus’.
F.2 Cognitive Profiles
To illustrate how a measurement layout identifies cognitive profiles, we explore a range of tasks in the Animal-AI Environment (Beyret et al., 2019; Voudouris et al., 2025). Full details of the environment can be found in Appendix B. we first select a simple navigation task where agents must navigate to a goal. In total, there are 69 selected navigation tasks and screenshots for two tasks are shown in Fig. 22. We include 68 agents submitted to the AAI Olympics competition (Crosby et al., 2020). The profiles of agents will be composed of navigation and perception skills and whether they have any bias towards goals occuring on the left vs. the right.
Following the analysis performed in (Burnell et al., 2022), we selected the set of meta-features $X$ to characterise task instances as rewardSize, rewardDistance, Xpos and rewardBehind. These represent the reward’s size, distance, and if it is left, right, in front, or behind the agent. Full details of the ranges of these variables and precise descriptions are presented in Appendix B, Table 10.
We then connect the task characterisation to the cognitive profile using a Measurement Layouts, that captures the interaction between demands and capabilities/bias. Fig. 24(a) presents the Measurement Layout we built. The metafeatures $X$ appear as observed (shaded, rectangular) nodes in the model, appearing within the plate, with 69 instances. The elements of the cognitive profile appear as external latent (unshaded) variables. The derived nodes lie at the intersection of the meta-features and the external cognitive profile nodes.
(a) Measurement layout for the Animal‑AI Olympics.
(b) Inferred cognitive profiles for the 68 AAIO agents. Error bars show s.d.; dot area encodes mean accuracy; colour encodes noiseLevel (red=high).
Figure 24:
Not visible in the figure are the linking functions associated with each arrow. Broadly, the relevant metafeature task demands combine to increase difficulty, and are opposed by relevant capabilities, yielding partial performance estimates in the forms of probabilities. In this task, because success requires both identidying the goal and navigating too it, the partial performances are multiplied to calculate the final probability of success.
Note that the ranges of these two abilities are given by the ranges of the task metafeatures, with navigationAbility in $[0,5.3]$ and with visualAbility in $[0,1.9]$ respectively, representing how distant and how big a reward—with their corresponding distance and size units—the agent could be measured as having in this task. This provides an intuitive property to the Measurement Layouts, in that the capabilities have meaningful units.
We detail all of the nodes and linking functions comprising this Measurement Layout in Table 11. The function $\sigma$ represents the standard logistic function, while $\omega$ represents a weighting:
$\omega[\alpha](b,c)=(1-\alpha)\cdot b+\alpha\cdot c$
The value $\nu$ represents a noise prior that represents a constant model set to 1 - average performance. Finally, $ScaledBeta$ represents a scaled beta distribution to an interval different from $[0,1]$. Note that a beta distribution has two parameters that need to be estimated. We describe our choices for building the measurement Layout in this way in Appendix F.
Name Type Range/Values Prior / Linking Function
rewardSize Meta-Feature $[0,1.9]$
rewardDistance Meta-Feature $[0,5.3]$
rewardBehind Meta-Feature $\{0,0.5,1\}$
XPos Meta-Feature $\{-1,0,1\}$
navigationAbility Capability $[0,5.3]$ $ScaledBeta(1,1,0,5.3)$
visualAbility Capability $[0,1.9]$ $ScaledBeta(1,1,0,1.9)$
rightleftBias Bias $\infty$$\infty$] $\mathcal{N}(0,1)$
noiseLevel Robustness $[0,1]$ $Uniform(0,1)$
rightleftEffect INON [1,1][-1,1] ${\sf rightleftBias}\times{\sf XPos}$
Table 11: Nodes in the AAIO Measurement Layout, including range/values, priors and linking function.
F.3 Inference and Analysis
Using the measurement layout and the 69 instances, we can use PyMC to infer the values of the cognitive profile for each of the 68 agents.
Fig. 24(b) summarises the inferred capabilities and noise levels. We see the two main capabilities are correlated for this population, however, we are also able to identify agents with weak navigation skill, but strong visual acuity (e.g.,cocel) and vice versa. We can also identify agents with strong biases (e.g., ‘Thursday’ and ‘KoozytHiperdyne’ with bias $1.14$ and 0.78-0.78). A full table with all of the inferred results for each agent is presented in Appendix F.
Where most evaluations end, with a results table of identified scores, Measurement Layouts are novel insofar as they can go on to predict agent performance on unseen tasks instances. This happens by the forwards inference process of the HBN, using the To evaluate this, we randomly held out 20% of the task instances from the backwards inference stage and evaluated the predicted probabilities of success on them. We evaluate these probabilities using a Brier score (Brier, 1950). That is:
$\textrm{Brier Score}(p,o)=\frac{1}{N}\sum_{i=1}^{N}(p_{i}-o_{i})^{2}$
Where $p_{i}$ and $o_{i}$ are the predicted probabilities of success and observed outcomes respectively. The Brier Score is essentially the mean square error applied to probabilities and binary events.
Due to the small size of the dataset, we repeated this process 15 times and took the mean Brier score for each agent. Figure 25 summarises these scores, which are lower than predictions based on aggregate performance for agents with success rates in the range $[0.2,0.7]$, exactly where there is variability to explain.
Figure 25: Brier score of model predictions against agent success rates for AAIO. The orange parabola is the expected value for the Brier score on aggregate predictions BS=s(1s)BS=s(1-s) for success rate $s$. The orange points are the observed Brier Scores for each agent using the aggregate prediction. These fall on the parabola as expected. The blue points are the Brier scores for our predictions using the Measurement Layout.
F.4 AAIO Patterns of Performance
To show the patterns of performance of several agents, we use a simple visualisation known as agent characteristic grids, as described in (Burnell et al., 2022). The plots place each feature as a dimension and the success rate represented in colours from green (success) to red (failure). Values are grouped into bins of appropriate width in the desired dimensions for analysis. Each cell also shows the number of observed instances in that bin. Fig. 26 shows some of the agents discussed in the paper, as well as some other agents that will serve us to illustrate a few points.
For instance, ‘ironbar’ has 99% accuracy (the grid is mostly green everywhere) for this set of tasks (like ‘Trrrr’, the winner of the competition) and is identified as saturating both capabilities in Fig. 24(b). The next one, ‘sparklemotion’, with 36% accuracy, is very non-conformant, with navigationAbility $0.40\pm 0.38$ and visualAbility $0.59\pm 0.45$, and high noise $0.31\pm 0.13$. Proportionally, navigationAbility is worse as the grid looks even less conformant on the distance of the reward than the reward size. The rightmost on the top row is ‘Juohmaru’, with 54% accuracy, which is a very conformant agent, with red concentrating on the top and right of the grid (the objectively most difficult instances). Its navigation ability is $2.50\pm 1.08$ and visual ability is $0.92\pm 0.44$, and medium noise: $0.26\pm 0.17$.
Appendix G Code and Datasets
@@
Our experiments did not require a large amount of compute, We ran all of our experiments on Google Colaboratory (Google, ) using the free tier of service. All corresponding code, datasets and raw results will be released on publication to ensure ease of replication.